# Gradient-flow SDEs have unique transient population dynamics

### v2

@@ 4 Identifiability results
5 Tri-level SB-refinement: nn-APPEX
We have shown in Section˜4 that the full SDE is identifiable from $N\geq 3$ distinct marginals, $\bigl(p(\cdot,t_{i})\bigr)_{i=1}^{N}$. Building on this, we propose Neural Network-based Alternating Projection Parameter Estimation from $X_{0}$ (nn-APPEX), an improved version of APPEX (Guan et al., 2024), which can learn the full SDE, beyond those with linear drift. nn-APPEX searches for the true parameters $(-\nabla\Psi,\sigma^{2})$ by alternating between trajectory inference, drift estimation, and diffusion estimation.
First, given a reference SDE $Q$, we solve a multi-marginal Schrödinger Bridge problem to infer a law on paths $P$. In particular, $P$ minimizes $\mathrm{KL}(P\|Q)$, while obeying the multi-marginal constraints, i.e. $P\in\Pi\bigl(p(\cdot,t_{i})_{i=1}^{N}\bigr)$. Second, from these inferred trajectories, we perform a maximum likelihood estimation (MLE) of the drift $\nabla\Psi$. For generality and tractability, we optimize over neural network parameters $\theta$. Third, we use MLE to estimate the diffusion $\sigma^{2}$. We then use the new drift and diffusion estimates to update the reference SDE $Q$ for the next iteration. Formally, at iteration $k$, nn-APPEX performs the steps
$P_{k}$ $=\arg\min_{P\in\Pi\bigl(p(\cdot,t_{i})_{i=1}^{N}\bigr)}\mathrm{KL}(P\|Q_{\shortminus\nabla\Psi_{\theta_{k-1}},\sigma^{2}_{k-1}}),$ (13)
$\theta_{k}$ $=\arg\min_{\theta}\mathrm{KL}(P_{k}\|Q_{\shortminus\nabla\Psi_{\theta},\sigma^{2}_{k-1}}),$ (14)
$\sigma^{2}_{k}$ $=\arg\min_{\sigma^{2}}\mathrm{KL}(P_{k}\|Q_{\shortminus\nabla\Psi_{\theta_{k}},\sigma^{2}}),$ (15)
and we iterate until convergence, or until a maximum number of iterations is reached. We emphasize that nn-APPEX does not require prior knowledge, since any gradient-flow SDE (1) can be used as the initial reference $Q_{-\nabla\Psi_{\theta_{0}},\sigma^{2}_{0}}$. While it is known that the $\mathrm{KL}$ divergence on the space $C([0,T],\mathbb{R}^{d})$ is infinite between SDEs with different diffusions (Shen et al., 2025), we note that given finite observation times, $\{t_{i}\}_{i=1}^{N}$, we may instead evaluate $\mathrm{KL}$ divergence on $C(\{t_{i}\}_{i=1}^{N},\mathbb{R}^{d})$, computed over the observed couplings $p_{t_{i},t_{i+1}}$. As shown in Guan et al. (2024, Section 4) and Appendix˜D, this objective is finite for gradient-flow SDEs (1), and thus enables principled tri-level optimization.
The SB problem (13) is equivalent to entropic OT, and can thus be solved with iterative proportional fitting (Peyré et al., 2019). For MLE parameter estimation on the inferred paths, we use the Euler-Maruyama approximation, and present derivations in Appendix˜E. In particular, the optimal drift parameters $\hat{\theta}$ are independent of $\sigma^{2}$, since minimizing the negative log-likelihood with respect to $\theta$ is equivalent to minimizing
$\ell(\theta)=\sum_{n=1}^{N}\sum_{k=0}^{T-1}\Big\|x_{(k+1)\Delta t}^{(n)}-x_{k\Delta t}^{(n)}+\nabla\Psi_{\theta}(x_{k\Delta t}^{(n)})\Delta t\Big\|_{2}^{2},$ (16)
where $N$ is the number of paths, $T$ is the number of time steps, and $\Delta t$ is the time step. We therefore estimate $\nabla\Psi_{\hat{\theta}}$ by fitting a neural network whose parameters minimize $\ell(\theta)$, as done in Shen et al. (2025). The diffusion MLE, $\hat{\sigma}^{2}$, is then derived from the quadratic variation, conditioned on the estimated drift $\nabla\Psi_{\hat{\theta}}$,
@@ 5 Tri-level SB-refinement: nn-APPEX
By alternating between these three procedures, nn-APPEX iteratively reduces the divergence between reconstructed paths, $P\in\Pi\left(p(\cdot,t_{i})_{i=1}^{N}\right)$, which obey the observed marginals, and paths from the estimated gradient-flow SDE with law $Q$:
@@
We note that $P$ need not be a gradient-flow SDE, and that $Q$ need not respect the marginal constraints. In fact, by ˜4.3, only one gradient-flow SDE $Q_{\shortminus\nabla\Psi,\sigma^{2}}$ obeys the marginal constraints, which implies
$\inf\limits_{\begin{subarray}{c}P\in\Pi\bigl(p(\cdot,t_{i})_{i=1}^{N}\bigr)\end{subarray}}\mathrm{KL}(P\|Q_{\shortminus\nabla\Psi,\sigma^{2}})=0\iff P=Q_{\shortminus\nabla\Psi,\sigma^{2}}.$
By decreasing the relative divergence via (18), nn-APPEX iteratively approaches the unique true solution. However, convergence to the unique minimizer is an open question. We discuss our algorithm’s convergence and runtime in Appendix˜D.
To highlight distinctions with existing SB methods, we note differences in algorithmic design, relative to nn-APPEX:
  • Waddington-OT (WOT) (Schiebinger et al., 2019) fixes the reference $Q$ to be a Brownian motion with diffusivity $\sigma^{2}$ and stops after (14).
  • SBIRR (Shen et al., 2025; Zhang, 2024) fixes $\sigma^{2}$ and iteratively performs steps (13) and (14).
  • APPEX (Guan et al., 2024) iteratively performs all three steps, but uses closed form MLE solutions, specific to linear additive noise SDEs, for the drift and diffusion updates.
@@
6 Experiments
In this section, we perform extensive experiments on simulated data to evaluate nn-APPEX against previous Schrödinger Bridge methods, WOT and SBIRR. The code repository is available on GitHub.
@@ 6.1 Experimental setup
To simulate gradient-flow SDEs, we fix the diffusivity $\sigma^{2}=0.2$ and consider five potentials that are commonly used as benchmarks in the literature (Terpin et al., 2024; Persiianov et al., 2025). Potentials were chosen to ensure the existence of a stationary distribution $p_{\mathrm{eq}}=\exp(-\frac{\Psi}{2\sigma^{2}})$. See Appendix˜F for the list of potentials. For our main experiments, we initialize the first marginal as a random Gaussian mixture model (GMM), due to their universal approximation properties (McLachlan and Rathnayake, 2014) and their applicability for clustering scRNA data into different cell types (Yu et al., 2021). We uniformly sample the number of components between $1$ and $10$, their means within $3,3]^{2}$, and their variances between $0.5$ and $1.0$. We then mirror the “three marginals” identifiability setting from ˜4.3, by sampling $N=2000$ points from the GMM initial distribution, then forward projecting these points using the Euler-Maruyama scheme for an additional $2$ time steps of size $\Delta t=0.01$. Experiments are repeated across $10$ seeds for each SDE to ensure replicability. Marginals for one seed are pictured for the SDE driven by the Oakley-O’Hagan potential in Fig.˜2(a)-(c).
@@ Methods.
To perform a systematic ablation that isolates the effects of iterative reference refinement and adaptive diffusion on inference, all methods are given the same data and initialization, and any shared subprocedure is implemented identically across methods. To emulate the realistic scenario where the practitioner can only estimate the true diffusivity $\sigma^{2}$ up to an order of magnitude (Guan et al., 2024), we sample the diffusivity prior as $\sigma^{2}_{\text{prior}}\sim\sigma^{2}\times 10^{\textrm{Unif}[-1,1]}$, for each of the $10$ seeds. This setting could be considered as favorable towards methods like WOT and SBIRR, since real-world diffusion estimates may in fact be off by two orders of magnitude (Forrow, 2024). For all methods, the SB step (13) is solved using up to $200$ iterations of a multi-marginal iterative proportional fitting procedure (IPFP) (Marino and Gerolin, 2020, Section 4.3), which rescales the per-time slice scalings until the marginal constraints are met up to an $L^{\infty}$ error of $10^{-5}$. $2000$ trajectories are then sampled from the inferred law on paths. Similarly, each method performs the drift MLE step (14) by fitting a $3$-layer multi-layer perceptron with silu activation, trained to minimize (16) via the Adam optimizer (epochs =500=500 and learning_rate $3\times 10^{-3}$) (Kingma and Ba, 2014). Only SBIRR and nn-APPEX are iterative, and we use $30$ iterations for each method.
@@
  • Cosine similarity: $\frac{\langle\nabla\Psi_{\hat{\theta}}(x),\hskip 2.84544pt\nabla\Psi(x)\rangle}{\|\nabla\Psi_{\hat{\theta}}(x)\|\|\nabla\Psi(x)\|}$.
@@ 6.2 Results
The results are aggregated over $10$ seeds for each gradient-flow SDE, and shown in Fig.˜3. We see that nn-APPEX yields the best and most robust estimates for the drift landscapes. On each of the five SDEs, its estimates have the lowest absolute error, the highest cosine similarity, and the lowest variance. While SBIRR significantly outperforms WOT, due to iterative drift refinement, its inability to update its diffusion prior prevents it from learning the drift field to the same fidelity as nn-APPEX. In particular, WOT and SBIRR regularly orient flow lines incorrectly in low potential zones, due to misspecified diffusivity (see Fig.˜2).
@@
We also replicate this experiment such that instead of a GMM initialization, the initial distribution is either uniform, $p_{0}\sim\textrm{Unif}[-4,4]^{2}$, as done in previous work (Terpin et al., 2024), or given by the stationary Gibbs distribution $p_{0}\sim p_{\mathrm{eq}}$. The results are summarized in Fig.˜4 and Fig.˜5 of the Supplement. For the uniform initialization, nn-APPEX continues to perform the best on all SDEs, and all methods perform relatively better, due to higher observability in the evaluation region $4,4]^{2}$. For the Gibbs initialization, all methods fail at inference, which is consistent with identifiability theory (˜4.1).
We also evaluate nn-APPEX against the state-of-the-art variational method JKOnet (Terpin et al., 2024) given GMM initializations. Results are plotted in Fig.˜6 and show that nn-APPEX achieves more accurate drift and diffusion estimates.
@@ 7 Conclusions
This work provides a call to action for researchers in ML and single-cell biology to move beyond the current paradigm of assuming known diffusion for marginals-based SDE inference, as we show that this is not only unnecessary, but also problematic in practice. Our theoretical contributions resolve a longstanding identifiability problem by proving that the gradient-flow drift and the diffusivity are jointly identifiable if and only if we observe transient marginals. This is significant for practical applications, since transience is a common and testable condition. We further expand the practicality of our result by proving that only three distinct marginals are needed for identifiability. To translate this theory into practice, we introduced the first Schrodinger Bridge-based method, capable of inferring arbitrary gradient-flow drift and diffusivity. Extensive numerical experiments demonstrate that our method, nn-APPEX, outperforms previous SB methods, in the realistic scenario where diffusivity is unknown.
@@
While our work contributes novel theory and methodology for SDE inference from marginals, we note several limitations, which point to important directions for future work. First, the gradient-flow SDE model (1) is prevalent in many fields, but it does not capture some important dynamics. In single-cell dynamics, one may expect rotational dynamics from non-conservative drift, due to genetic feedback loops. As another example, in hydrology, material heterogeneity requires modified models for drift and diffusion. Second, our theory was proven given exact observation of the marginals $p(\cdot,t)$. Rigorous asymptotic theory quantifying identifiability with finite-samples and observational noise would be useful to the community. Finally, there is still a gap in showing optimality of joint inference methods. While experiments show positive results, proving that nn-APPEX converges to the true parameters under basic conditions would mark a major advance.

### v3

@@ 4 NN-APPEX: THREE-STAGE SB REFINEMENT
The SB problem (13) is equivalent to entropic optimal transport with an adjusted cost, and can thus be solved with a variety of iterative proportional fitting algorithms, such as Sinkhorn’s algorithm (Peyré et al., 2019). By default, we use the multi-marginal Sinkhorn algorithm (Marino and Gerolin, 2020). For MLE parameter estimation on the inferred paths, we use the Euler-Maruyama approximation, and present derivations in Appendix˜E. In particular, the optimal drift parameters $\hat{\theta}$ are independent of $\sigma^{2}$, since minimizing the negative log-likelihood with respect to $\theta$ is equivalent to minimizing
@@
where $M$ is the number of paths, $N$ is the number of time steps, and $\Delta t$ is the time step. We therefore estimate $\nabla\Psi_{\hat{\theta}}$ by fitting a neural network whose parameters minimize $\ell(\theta)$, as done in Shen et al. (2025). The diffusion MLE, $\hat{\sigma}^{2}$, is then derived from the quadratic variation, conditioned on the estimated drift $\nabla\Psi_{\hat{\theta}}$,
[…]
Algorithm 1 $n$-APPEX
1:Input: Observed marginals $\hat{p}_{t_{i}}$, $i=0,\ldots,N-1$, number of iterations $K$, time step $\Delta t$
2:Initialize: $\nabla\hat{\Psi}\leftarrow 0$,  $\hat{\sigma}^{2}\leftarrow 1$
3:for $k=1,\ldots,K$ do
4:  $\{\mathcal{K}_{i}(x,y)\}_{i=0}^{N-2}\leftarrow\exp\!\left(-\frac{\|y-x+\nabla\hat{\Psi}(x)\,\Delta t\|^{2}}{2\hat{\sigma}^{2}\Delta t}\right)$
5:  $\{\pi_{i}\}_{i=0}^{N-2}\leftarrow\textsc{MultiSinkhorn}\!\left(\{\hat{p}_{t_{i}}\},\,\{\mathcal{K}_{i}\}\right)$
6:  $\{\hat{\tau}^{(s)}\}_{s=1}^{S}\leftarrow\textsc{SamplePaths}\!\left(\{\pi_{i}\}\right)$
7:  $\nabla\hat{\Psi}\leftarrow\textsc{NNTrain}\!\left(\{\hat{\tau}^{(s)}\}\right)$
[…]
9:end for
10:return $\nabla\hat{\Psi},\;\hat{\sigma}^{2}$
By alternating between these three procedures, nn-APPEX iteratively reduces the divergence between reconstructed paths, $P\in\Pi\left(p(\cdot,t_{i})_{i=0}^{N-1}\right)$, which obey the observed marginals, and paths from the estimated gradient-flow SDE with law $Q$:
@@
By decreasing the relative divergence, nn-APPEX iteratively approaches the unique true solution, though it is unclear whether the algorithm can stagnate at a different set of parameters. We conjecture that convergence to the true solution holds given infinite data, and leave the proof for future work.
To highlight distinctions with existing SB methods, we note relative differences in algorithmic design:
  • Waddington-OT (WOT) (Schiebinger et al., 2019) fixes the reference $Q$ to be a Brownian motion with diffusivity $\sigma^{2}$ and stops after (14), i.e., fix $K=1$ in Algorithm˜1.
  • SBIRR (Shen et al., 2025; Zhang, 2024) fixes $\sigma^{2}$ and iteratively performs steps (13) and (14), i.e., exclude line 8 in Algorithm˜1.
  • APPEX (Guan et al., 2024) iteratively performs all three steps, but only considers linear drift, i.e., replace line 7 in Algorithm˜1 with a closed form MLE for linear SDEs.
@@
5 SIMULATED EXPERIMENTS
In this section, we perform extensive experiments on simulated data to evaluate nn-APPEX against previous Schrödinger Bridge methods, WOT and SBIRR. The code repository is available on GitHub: https://github.com/guanton/identifying-gradient-flow-sdes.
@@ 5.1 Experimental Setup
To simulate gradient-flow SDEs, we fix the diffusivity $\sigma^{2}=0.2$ and consider five potentials that are commonly used as benchmarks in the literature (Terpin et al., 2024; Persiianov et al., 2025). Potentials were chosen to ensure the existence of a stationary distribution $p_{\mathrm{eq}}=\frac{1}{Z}\exp(-\frac{\Psi}{2\sigma^{2}})$. See Appendix˜F for the list of potentials. For our main experiments, we initialize the first marginal as a random Gaussian mixture model (GMM), due to their universal approximation properties (McLachlan and Rathnayake, 2014) and their applicability for clustering scRNA data into different cell types (Yu et al., 2021). We uniformly sample the number of components between $1$ and $10$, their means within $3,3]^{2}$, and their variances between $0.5$ and $1.0$. We then mirror the “three marginals” identifiability setting from ˜3.3, by sampling $N=2000$ points from the GMM initial distribution, and then forward simulating these points using the Euler-Maruyama scheme for an additional $2$ time steps of size $\Delta t=0.01$. Experiments are repeated across $10$ seeds for each SDE to ensure replicability. Marginals for one seed are pictured for the SDE driven by the Oakley-O’Hagan potential in Figure˜2(a)-(c).
@@ Methods.
To perform a systematic ablation that isolates the effects of iterative reference refinement and adaptive diffusion on inference, all methods are given the same data and initialization, and any shared subprocedure is implemented identically across methods. To emulate the realistic scenario where the practitioner can only estimate the true diffusivity $\sigma^{2}$ up to an order of magnitude (Guan et al., 2024), we sample the diffusivity prior as $\sigma^{2}_{\text{prior}}\sim\sigma^{2}\times 10^{\textrm{Unif}[-1,1]}$, for each of the $10$ seeds. This setting could be considered favourable towards methods like WOT and SBIRR, since real-world diffusion estimates may in fact be off by two orders of magnitude (Forrow, 2024). For all methods, the SB step (13) is solved using up to $200$ iterations of a multi-marginal iterative proportional fitting procedure (IPFP) (Marino and Gerolin, 2020, Section 4.3), which rescales the per-time slice scalings until the marginal constraints are met up to an $L^{\infty}$ error of $10^{-5}$. $2000$ trajectories are then sampled from the inferred law on paths. Each method performs the drift MLE step (14) by fitting a $2$-layer (128 neurons per layer) multi-layer perceptron with SiLU activation, trained to minimize (16) via the Adam optimizer (epochs =500=500 and learning_rate $3\times 10^{-3}$) (Kingma and Ba, 2014). Only SBIRR and nn-APPEX are iterative, and we use $30$ iterations for each method.
@@
  • Cosine similarity: $\frac{\langle\nabla\Psi_{\hat{\theta}}(x),\hskip 2.84544pt\nabla\Psi(x)\rangle}{\|\nabla\Psi_{\hat{\theta}}(x)\|\|\nabla\Psi(x)\|}$
@@ 5.2 Results
The results are aggregated over $10$ seeds for each gradient-flow SDE, and shown in Figure˜3. We see that nn-APPEX yields the best and most robust estimates for the drift landscapes. On each of the five SDEs, its estimates have the lowest absolute error, the highest cosine similarity, and the lowest variance. While SBIRR significantly outperforms WOT, due to iterative drift refinement, its inability to update its diffusion prior prevents it from learning the drift field to the same fidelity as nn-APPEX (see also Table˜4 and Table˜5 in the Appendix). In particular, WOT and SBIRR regularly orient flow lines incorrectly in low potential zones, due to misspecified diffusivity (see Figure˜2 and Table˜4).
@@
We also replicate the experiment such that the initial distribution is either uniform, $p_{0}\sim\textrm{Unif}[-4,4]^{2}$, as done in previous work (Terpin et al., 2024), or given by the stationary Gibbs distribution $p_{0}\sim p_{\mathrm{eq}}$. The results are summarized in Figure˜4 and Figure˜5. For the uniform initialization, nn-APPEX continues to perform the best on all SDEs, and all methods perform relatively better, due to higher observability in the evaluation region $4,4]^{2}$. For the Gibbs initialization, all methods fail at inference, which is consistent with identifiability theory (˜3.1). Finally, we evaluate nn-APPEX against the state-of-the-art variational method JKOnet (Terpin et al., 2024) given GMM initializations. Results are plotted in Figure˜6 and show that nn-APPEX achieves more accurate drift and diffusion estimates.
6 BIOLOGICAL EXPERIMENTS
We now test nn-APPEX’s ability to perform trajectory inference on real biological data. As done in Shen et al. (2025), we use single-cell data from human embryonic stem cells (hESC) (Chu et al., 2016).
6.1 Experimental Setup
For consistency, we use the same data preprocessing, method implementation, and performance evaluation as Shen et al. (2025). Our code builds on the SBIRR public repository, such that the only modifications were to randomize the initial diffusivity, and to add the diffusion update step to implement nn-APPEX.
Dataset.
The hESC dataset (Chu et al., 2016) comprises $5$ time marginals (0h, 1212h, 2424h, 3636h, 7272h), and the observed number of cells per marginal are: $92$, $102$, $66$, $172$, $138$. 55 PCA components are used.
Method implementation.
For this experiment, the SB solver for each method is implemented with the forward-backward iterative proportional maximum likelihood algorithm (Vargas et al., 2021). SBIRR and nn-APPEX both use the same gradient field MLE for re-estimating the reference drift (see Shen et al. (2025, Section D.6.2) for details), and we add the closed form diffusivity MLE (15) for re-estimating the reference diffusion for nn-APPEX. We also implement a time-varying version of nn-APPEX, which computes (15) between each pair of marginals to estimate time-varying diffusivity. All iterative methods are run for $10$ iterations.
Evaluation.
We train each method on the first (0h), third (24h), and fifth (72h) marginals, and then use the learned dynamics to predict the second (12h) and fourth (36h) marginals. We then evaluate performance using the earth mover’s distance between the estimated and true holdout marginals. We average results over $10$ random seeds, and for each seed, we sample a random initial diffusivity within an order of magnitude of $\sigma^{2}=0.1$, the default considered in Shen et al. (2025).
6.2 Results
We report results in Table 1. We observe that SBIRR’s iterative drift refinement enables better performance compared to the simplest method WOT. Adding the diffusion update then provides another small but noticeable improvement, such that nn-APPEX (with time-varying diffusivity) achieves the best overall performance on both holdout marginals.
This experiment enhances our point that diffusion should be learned from the data, rather than assumed beforehand. Indeed, for real data, the diffusion is unknown, and it is often not obvious how it should be specified. Methods that refine both drift and diffusion also predict unseen marginals with higher accuracy, which suggests that they learn a better model.
Table 1: The average earth mover’s distance ($\pm$ standard error) across each method and each hold out time on the hESC dataset. The best results are displayed in bold and the second best results are underlined.
Method Time Avg. EMD
WOT $t_{2}$ 0.803 $\pm$ 0.039
$t_{4}$ 1.492 $\pm$ 0.081
SBIRR $t_{2}$ 0.694 $\pm$ 0.021
$t_{4}$ 1.473 $\pm$ 0.033
$n$-APPEX $t_{2}$ 0.683 $\pm$ 0.021
$t_{4}$ 1.470 $\pm$ 0.040
nn-APPEX (time-varying) $t_{2}$ 0.678 $\pm$ 0.022
$t_{4}$ 1.454 $\pm$ 0.033
7 RELATED WORKS
Our work contributes novel theory and methodology for identifying the true SDE from observed marginals. We review relevant literature below.
Identifiability theory for population dynamics.
A popular approach for marginals-based inference is to reconstruct trajectories with entropic optimal transport, which is equivalent to solving a Schrödinger Bridge (SB) problem (Peyré et al., 2019, Prop. 4.2). The first such method was Waddington-OT (WOT) (Schiebinger et al., 2019), which performs trajectory inference by solving a single SB problem with respect to a Brownian motion reference with a prescribed diffusivity $\sigma^{2}$. WOT spawned many variations, which incorporate additional information, like cell lineages (Forrow and Schiebinger, 2021; Ventre et al., 2023), or improve robustness given limited samples (Zhang et al., 2021; Lavenant et al., 2024; Chizat et al., 2022). In the last year, SB methods were improved with the introduction of iterative reference refinement, i.e. SBIRR (Shen et al., 2025; Zhang, 2024). Rather than fixing Brownian motion as a reference, SBIRR fixes the diffusivity $\sigma^{2}$ and iteratively re-estimates the drift of the reference process. In contrast to SBIRR, the method that we introduce in this work, nn-APPEX, additionally re-estimates the diffusivity at each iteration. nn-APPEX builds upon the three-stage iterative scheme of APPEX (Guan et al., 2024).
Other inference methods for population dynamics.
In addition to SB methods, variational energy-based methods (Hashimoto et al., 2016; Bunne et al., 2022; Neklyudov et al., 2023; Terpin et al., 2024; Persiianov et al., 2025) have been developed for SDE inference from population dynamics. These methods leverage the fact that the marginals of potential-driven SDEs minimize a corresponding energy, and use discrete numerical schemes to approximate the underlying parameters. Of these, only JKOnet (Terpin et al., 2024) jointly estimates drift and diffusion. We note that JKOnet and nn-APPEX consider distinct optimization approaches (the former leverages the JKO scheme (Jordan et al., 1998) with a joint variational step, whereas nn-APPEX performs three-stage SB refinement) and we compare their performance in Appendix˜G. We also note that Terpin et al. (2024) did not study identifiability theory, and therefore lacked guarantees for principled inference. Recent work presents another framework for joint inference via maximum mean discrepancy minimization (Berlinghieri et al., 2025).
@@ 8 CONCLUSIONS
This work provides a call to action for researchers in ML and single-cell biology to move beyond the current paradigm of assuming known diffusion for marginals-based SDE inference, as we show that this is not only unnecessary, but also problematic in practice. Our theoretical contributions resolve a longstanding identifiability problem by proving that the gradient-flow drift and the diffusivity are jointly identifiable if and only if we observe transient marginals. This is significant for practical applications, since transience is a common and verifiable condition. We further expand the practicality of our result by proving that only three distinct marginals are needed for identifiability. To translate this theory into practice, we introduced the first Schrödinger Bridge-based method capable of inferring arbitrary gradient-flow drift and diffusivity. Extensive experiments demonstrate that our method, nn-APPEX, outperforms previous SB methods, in the realistic scenario where diffusivity is unknown.
@@
While our work contributes novel theory and methodology for SDE inference from marginals, we note several limitations, which point to important directions for future work. First, the gradient-flow SDE model (1) is prevalent in many fields, but it does not capture some important dynamics. In single-cell dynamics, one may expect rotational dynamics from non-conservative drift, due to genetic feedback loops. As another example, in hydrology, material heterogeneity requires modified models for drift and diffusion. Second, our theory was proven given exact observation of the marginals $p(\cdot,t)$. Rigorous asymptotic theory quantifying identifiability with finite samples and observational noise would be useful to the community. Finally, there is still a gap in showing optimality of joint inference methods. While experiments show positive results, proving that nn-APPEX converges to the true parameters under basic conditions would mark a major advance.